OpenStudy (studyboiiiiiii):

Solve for x: -3|2x + 6| = -12 A. x=1, and x=5 B. x=-1, x=-5 C. x==9 and x=3 D. No solution

OpenStudy (studyboiiiiiii):

|2x + 6| =-9 ?

563blackghost (563blackghost):

So so sorry I meant you had to divide by 3 from each side....

563blackghost (563blackghost):

by -3 that is.... \(\huge{\frac{-3}{-3}|2x+6|=\frac{-12}{-3}}\) What would that equal?

OpenStudy (kevin):

I will make it easier... -3|2x + 6| = -12 Divide both sides by 3, It will be like this : |2x + 6| = 4 You have 2 possibility answers: i (2x + 6) = 4 ii -(2x + 6) = 4 Find x value on first and second equation above